OpenStudy (anonymous):

At noon, ship a is 50 nautical miles due west of ship b. Ship a is sailing west at 16 knots and ship b is sailing north at 15 knots. How fast in knots is the distance between the ships changing at 4 pm?

OpenStudy (anonymous):

Let s be the total distance between them, x horizontal distance and y the vertical distance \[ s^2=x^2+y^2 \]By chain rule, we get:\[ 2ss'=2xx'+2yy' \]

OpenStudy (anonymous):

We want to know \(s'\).

OpenStudy (anonymous):

You got to figure out \(x, y, z\) at 4 pm. Though we already know \(x'=16, y'=15\).

OpenStudy (anonymous):

I mean \(s\) not \(z\).

OpenStudy (anonymous):

I am not certain I know how to put it together to find x,y, and s...

OpenStudy (anonymous):

Well, they tell you starting at noon, so at 4pm, 4 hours have passed. We have \(x = 50+4(16)\) and \(y=4(15)\) and \(z = \sqrt{x^2+y^2}\)

OpenStudy (anonymous):

Again, when I say z or s I mean the same thing, sorry.

OpenStudy (anonymous):

So i just needed to draw a right triangle and solve for the hypotenuse (spellingsorry)? Take the derivative of that equation plug in t=4 and get my answer... Took me 4 hours to figure it out.. Thank you!!!